OpenStudy (anonymous):

square root of 3 sec x=2

OpenStudy (solomonzelman):

\[\sqrt{3}Secx=2~~~~~same~~as~~~\sqrt{3} \times Secx=2\]\[\sqrt{3} \times Secx \color{red} {\div \sqrt{3}}=2 \color{red} {\div \sqrt{3}}~~~~~->~~~~~\color{blue} {Secx=\frac{2}{\sqrt{3}}}\] \[\color{blue} {Secx=\frac{2\color{red} {\times \sqrt{3}}}{\sqrt{3\color{red} {\times \sqrt{3}}}}}\color{blue}{=\frac{2\sqrt{3}}{3}}\]

OpenStudy (solomonzelman):

Now use the trigonometric table to find x.